# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ B.2.2 Out-of-distribution
In this section we investigate the how robustly a neural network has learned visual concepts: how well can they recognize objects under different, more challenging settings? This is an important question for many applications, and gives an idea both of the generalization properties, as well as to the practical applicability of models. To investigate robustness, we look at six different out-of-distribution testsets for the popular ImageNet dataset in Fig. 13. ImageNet ReAL [Beyer et al., 2020] relabels the dataset to fix inaccuracies/oversights in the original data. ImageNet V2 [Recht et al., 2019] is a new, updated test-set for the original dataset. ImageNet-A [Hendrycks et al., 2021b] consists of adversarial examples, whose aim is to confuse/trick modern systems. ImageNet-C [Hendrycks and Dietterich, 2019] tests how well models can deal with corrupted inputs, e.g. very noisy or blurry images. Finally, ImageNet-R [Hendrycks et al., 2021a] shows more abstract renditions of actual objects, for example scetches, sculptures or drawings.
@@
As we see in the results, AligNet finetuning has a positive impact on the models under investigation, with most datasets showing slight improvements. On average over all models and across all datasets, accuracy went up by 2.6 percentage points, which on Imagenet-like datasets is considered a large performance boost [Beyer et al., 2020]. As can be expected, the models that were pre-trained on imagenet already showed the smallest improvements. Even on ImageNet ReAL, which can be considered a more reliable measure of model improvement than the original Imagenet, on average the models improved by 1.5%. The clearest advantage of AligNet can be seen for adversarial examples where AligNet-finetuned models exhibit much better behavior than their baseline counterparts: the average improvement was a jump of 5.6 percentage points. For corruptions and renditions, the situation is mixed: while on average the results still show an improvement of 1.8% on ImageNet C and 1.4% on ImageNet R, the SigLIP models slightly deteriorated. Overall, this fits with our understanding of what AligNet is meant to achieve: it allows models to learn clearer concepts that generalize better, but it is not improve performance in situations where image corruption or artistic renditions are the cause of misclassifications.
Figure 13: Out of distribution results on variations of ImageNet.
B.2.3 Ablations
@@
To investigate if AligNet finetuning affects models of different capacity/size in a similar way, we compared the performance of the ViT-B model (87M parameters) with the performances of a smaller ViT-S (22M parameters) and a larger ViT-L (305M parameters) model respectively. The results in Fig. 12 and Fig. 13 indicate that the positive impact of AligNet fine-tuning increases with model scale. For example, the average 10-shot performance improved by $2.4\%$ points for ViT-S, by $6.3\%$ points for ViT-B and by $8.3\%$ points for ViT-L. We observe similar results for all datasets we considered. Thus, we hypothesize that a larger model capacity makes it easier for a model to integrate the additional global coarse-grained information provided by AligNet’s human-like similarity judgments.
@@
We also wanted to see how AligNet scales with the amount of available training data. We individually varied the number of ImageNet images we used to construct our dataset over 3 orders of magnitude, and the number of triplets that we sampled across 4 orders of magnitude. We then looked at the 1-shot results across all the few-shot datasets described in Sec. B.2.1 (Outcomes for 10-shot were not qualitatively different). The results in Fig. 14a show that AligNet performance only increases slightly as the number of available base images increases. However, as we can see from Fig. 14b, the number of generated triplets does influence the outcome, though we hit diminishing returns eventually. We do not expect to be able to increase AligNet performance much further be simply increasing the amount of available data.
Figure 15: A: Odd-one-out accuracies on the THINGS dataset for different models in different settings. This figure is similar to (an extended version of) the THINGS plot in Fig. 2A. “AligNet” and “Base Model” refer to the pretrained student model, without and without finetuning on AligNet respectively. UnAlignet refers to a student finetuned on a variant of the AligNet dataset with an unaligned teacher. B: Comparing 1-shot accuracy on 11 datasets after fine-tuning on UnaligNet ($x$-axis) with fine-tuning on AligNet ($y$-axis). Most points lie slightly above the diagonal indicating that alignment of the teacher leads to a small but significant improvement in 1-shot performance of the student. This ablation shows the importance of aligning the teacher model for getting aligned student models.
@@
An important part of the AligNet pipeline is to align the teacher model to human similarity judgments using a specially trained affine transformation. Here we investigate the effect of this step, by generating a variant of AligNet without this transformation that we call UnAligNet. For this dataset we do not use any human odd-one-out data, so finetuning on UnAligNet is equivalent to distilling the teacher model using our triplet sampling procedure. Fig. 15A compares the odd-one-out accuracy of various student models when fine-tuning on AligNet vs on UnaligNet. As expected, we find that using an unaligned teacher leads to substantially worse alignment of the student models. Though it should be noted that some models (especially the ViTs trained on ImageNet) can already profit even from an unaligned teacher.
It is perhaps unsurprising that an aligned teacher improves the alignment of the student models. A more interesting question is, whether an aligned teacher also improves the students performance on other tasks such as few-shot classification. Fig. 15B thus compares 1-shot performance after training on UnaligNet ($x$-axis) with performance after training on AligNet ($y$-axis). The first thing to note here, is that the differences between AligNet and UnaligNet are much smaller than the differences between AligNet and the Base Model (cf. Fig. 12). This shows that the few-shot improvements are in large part due to the teacher model and our distillation method. But importantly, the vast majority of remaining improvements are positive, which means that using an aligned teacher gives a consistent advantage over an unaligned teacher.
@@
Fig. 16 shows the 1-shot accuracy of student models when comparing different sampling strategies. In all three panels, the $y$-axis corresponds to the performance of a model after AligNet-fine-tuning. The left panel, again shows how this compares to the performance of the (not fine-tuned) base model ($x$-axis). The middle panel compares shows how AligNet performance compares to fine-tuning on random-triplets ($x$-axis). Finally, the right panel shows comparison between AligNet and fine-tuning on triplets sampled from class-borders. Notice that the difference in the middle panel between random triplet sampling is consistently and substantially worse than cluster-based sampling. This clearly highlights the importance of the triplet sampling for down-stream performance. The panel on the right, on the other hand, shows that there is virtually no difference between cluster-border sampling and class-border sampling. So when label information is available, that can effectively be used for sampling triplets, but even if there is not, the clustering based sampling will work just as well.
@@
We have stressed that fine-tuning models to better align them with human similarity judgments depends on having sufficiently large datasets, which do not exist yet. This argument has motivated our multi-stage approach of first linearly aligning a teacher model and then using this as a surrogate model to generate a large corpus of human-like triplet odd-one-out choices. Yet, if we look at Fig. 14, it seems that decreasing the number of images or the number of triplets by up to two orders of magnitude still retains most of the benefit on downstream task performance (in terms of few-shot accuracy). This raises the question whether our pipeline as a whole might be unnecessary, given that the THINGS dataset [Hebart et al., 2023] already contains 4.70 million human similarity judgments collected via online crowd- sourcing for 1854 object images.
Therefore, here we compare soft-alignment with using the THINGS odd-one-out triplets directly for fine-tuning student models. Fig. 17 shows that fine-tuning directly on THINGS is consistently worse compared to training on AligNet in terms of ImageNet 1-shot accuracy ($x$-axis). Importantly, it appears that in almost all cases the performance of the model actually decreases below that of the base model. It is further interesting to note that even in terms of THINGS odd-one-out accuracy, the AligNet trained models perform better than the any other model. This is somewhat surprising, because it means that there clearly exists an overfitting problem when training on the THINGS data (even with strong $\ell_{2}$-regularization), which seems to mostly disappear when using AligNet for fine-tuning. One reason for this could be the differences in the sampling procedure (cluster-based vs. random sampling) and the loss function (soft-choices vs. hard-choices) between AligNet and THINGS.
@@
Uncertainty Distillation vs. gLocal. Here, we compare the Uncertainty Distillation (UD) transformation that we introduced in Appx. A.1.1 against gLocal [cf. Muttenthaler et al., 2023b]. The learning objectives of the two linear transformations are highly similar and the goal is the same: the representations of a pretrained vision foundation model are linearly transformed into a global coarse-grained human object similarity space using the THINGS dataset [Hebart et al., 2020] while the nearest neighbor structure of the model’s representation space is preserved. However, the difference between gLocal and UD is that UD directly injects the human uncertainty estimates for the triplet odd-one-out choices via approximate Bayesian inference [cf. Muttenthaler et al., 2022] into a model’s representation space and, thus, uses soft rather than hard triplet choices (see Appx. A.1.1 for details). We use the Levels datasets (see Appx. A.3) for evaluating the UD and gLocal transformations, comparing the same set of models that were evaluated in Muttenthaler et al. [2023b]. The gLocal transformations for those models are publicly available via the Python library thingsvision [Muttenthaler and Hebart, 2021] which we used for this analysis. In Fig. 18 we see that the UD transformation is either equally accurate or better in aligning the model’s representations with the human similarity judgments, and predicting their disagreement levels and aggregated RTs respectively in the different abstraction level compared to gLocal. As before, human RTs are measured in log-space and model uncertainties are measured as discrete Shannon entropy over the triplet odd-one-out choices. Note that UD is computationally more efficient than gLocal because it does not require an extra set of images—in addition to the THINGS data—to evaluate the preservation of a representation’s nearest neighbor structure which the multi-objective of gLocal necessitates [cf. Muttenthaler et al., 2023b].

### v4

@@ B.2.2 Out-of-distribution
Figure 13: Out of distribution results on several popular variations of ImageNet. Accuracy of a linear readout head trained on top of the (frozen) base models is plotted on the $x$-axis, against the corresponding performance a linear readout head trained on top of the (frozen) AligNet-finetuned models on the $y$-axis. Every symbol above the diagonal indicates improved performance (through AligNet), while symbols below the diagonal indicate a performance degradation. Note that for better readability the range of the plot is zoomed in on the pre-trained models. The model trained from scratch ($\times$) is thus often out of bounds due to its substantially lower performance, and its accuracy before and after finetuning is instead written out in a small inset.
@@
In this section we investigate the how robustly a neural network has learned visual concepts: how well can they recognize objects under different, more challenging settings? This is an important question for many applications, and gives an idea both of the generalization properties, as well as to the practical applicability of models. To investigate robustness, we look at six different out-of-distribution testsets for the popular ImageNet dataset in Fig.˜13. ImageNet ReAL [5] relabels the dataset to fix inaccuracies/oversights in the original data. ImageNet V2 [68] is a new, updated test-set for the original dataset. ImageNet-A [32] consists of adversarial examples, whose aim is to confuse/trick modern systems. ImageNet-C [30] tests how well models can deal with corrupted inputs, e.g. very noisy or blurry images. Finally, ImageNet-R [31] shows more abstract renditions of actual objects, for example scetches, sculptures or drawings.
@@
As we see in the results, AligNet finetuning has a positive impact on the models under investigation, with most datasets showing slight improvements. On average over all models and across all datasets, accuracy went up by 2.6 percentage points, which on Imagenet-like datasets is considered a large performance boost [5]. As can be expected, the models that were pre-trained on imagenet already showed the smallest improvements. Even on ImageNet ReAL, which can be considered a more reliable measure of model improvement than the original Imagenet, on average the models improved by 1.5%. The clearest advantage of AligNet can be seen for adversarial examples where AligNet-finetuned models exhibit much better behavior than their baseline counterparts: the average improvement was a jump of 5.6 percentage points. For corruptions and renditions, the situation is mixed: while on average the results still show an improvement of 1.8% on ImageNet C and 1.4% on ImageNet R, the SigLIP models slightly deteriorated. Overall, this fits with our understanding of what AligNet is meant to achieve: it allows models to learn clearer concepts that generalize better, but it is not improve performance in situations where image corruption or artistic renditions are the cause of misclassifications.
In their comprehensive survey on the robustness of computer vision models, Liu et al. [49] compare state-of-the-art training methods for improving the robustness of Vision Transformers (“ViTs”), namely Adversarial Training (“AT”), Masked Autoencoder (“MAE”), and Contrastive Language-Image Pretraining (“CLIP”). These methods have been shown to substantially improve robustness upon vanilla supervised or pure image self-supervised training. Similarly to AT, our method can be employed in addition to the other four methods evaluated in that survey. In our work we align SSL models similar to MAE, such as DINO and DINOv2, and models pretrained using CLIP (we use SigLIP which is a variant of CLIP). We find that soft-alignment performs comparably to AT (e.g., ImageNet-A, ImageNet-R). On average, it is neither significantly worse nor significantly better than AT and substantially improves upon the non-aligned base model representations (see Tab. 6). Since ViT-B/16 is the current defacto standard architecture in computer vision, we use it as the basis for all of our soft-aligned models and compare against the numbers for ViT-B reported in Liu et al. [49] (see Tab. 6). We note that while our method is not primarily targeted toward optimizing OOD robustness (or has an adversarial training component), improved OOD robustness appears to be a useful side-effect that emerges from transforming the model representations into a more human-aligned and interpretable space via our method.
IID Real-world OOD Synthesized OOD
Backbone (Pre-)Training IN-Val IN-21K IN-V2 IN-Real ON IN-A IN-R IN-V IN-C
ViT-B/16 Vanilla 75.7 91.7 61.6 80.9 20.8 11.3 32.8 24.2 34.3
MAE 83.6 90.5 73.1 88.1 37.4 37.4 49.8 36 49.4
CLIP 68.4 69.8 61.9 75.1 45 50 77.7 39.4 29.3
AT 73.4 86.7 60.4 80.6 19.2 8.8 50.7 20.2 36.6
Vanilla + AligNet 77.6 65.4 83.6 15.2 40.2 72.3
SigLIP + AligNet 82.8 73.3 87 37 81.4 74.8
Table 6: Comparison of state-of-the-art (pre-)training methods for improving downstream task performance of computer vision models [cf. 49] and soft-aligned models (i.e., models fine-tuned on AligNet) for various ImageNet-based evaluation datasets, testing a model’s performance either on i.i.d. data or its OOD robustness. Boldface indicates the best performance for a dataset.
B.2.3 Distribution shift
Figure 14: Additional results for all student models on the four Breeds datasets: “entity13”, “entity30”, “living17”, and “nonliving26”. Student models are sorted according to their objective function (supervised, captioning, self-supervised, image/text). The randomly initialized ViT-B has the poorest base performance but benefits the most (leftmost column in each panel). Conversely, SigLIP-So400m—the teacher model—has the strongest base performance and benefits the least but consistently (rightmost column in each panel). Vertical arrows indicate improvements in performance compared to the base performance.
To measure whether the global category structure induced by fine-tuning on AligNet helps to alleviate model problems with distribution shifts between the training and the test set, we evaluate our models on the BREEDS benchmarks [74]. These were specifically designed to test generalization under input distribution shifts, by constructing datasets where the samples in the training and test sets are sampled from different subpopulations. For example, the training set for the “dog” class contains some selected dog breeds in the training set (e.g., Dalmatians and Schnautzers), whereas a disjoint set of dog breeds is used for the test set (e.g.. Poodles and German Shepards).
In this section we expand on the results presented in Sec. 2.3 in the main text and evaluate all of our AligNet fine-tuned student models (see Sec. A.1.5 for details) on BREEDS. For each model, we train a linear probe on the four BREEDS training sets—“entity13”, “entity30”, “living17”, and “nonliving26”—, and evaluate that probe on the corresponding test sets.
AligNet fine-tuning consistently improves performance for all model types across all of the BREEDs benchmarks. We find that the poorer the baseline performance of a student model is, the more that model benefits from fine-tuning on AligNet. For example, the baseline performance of the randomly initialized ViT-B model (which we refer to as “Scratch”) is by far the worst but after AligNet fine-tuning that model performs as well as the pretrained student models for “entity13” and “entity30”. For the “living17” and “nonliving26” datasets its performance is slightly worse (but much less than before fine-tuning) than that of the other models, which may be due to the missing pretraining step. Conversely, SigLIP-So400m—the teacher model—has the strongest base performance and thus benefits the least from finetuning. It is striking, however, that even the teacher model consistently improves on BREEDS although that is the model that we used to generate AligNet. For “entity30” some of the student models even outperform the teacher model. This is suggestive evidence that the human part of the AligNet data plays a larger role for these kind of analyses than the teacher model part.
Another interesting observation is that larger supervised ViT models have worse base performance than smaller supervised ViT models but benefit notably more from AligNet fine-tuning than the smaller models. Before AligNet fine-tuning the order of the models according to their base performance is ViT-S, ViT-B, ViT-L across all four datasets. This order is reversed after AligNet fine-tuning for all four datasets. Thus, similarly to the human alignment results presented in Fig.˜7, there seems to exist a small scaling effect where larger models benefit more from AligNet fine-tuning, at least for the models that are not the teacher itself. We show all model results for the BREEDS dataset in Fig.˜14.
B.2.4 Additional Experimental Results
CLIP.
Figure 15: Results for 1-shot and 10-shot linear probing on a number of common few-shot evaluation datasets for CLIP, SigLIP, and SigLIP2.
In the main paper we have focused on SigLIP [85] as the representative model for image/text contrastive models because it is a very strong model that significantly improved upon the more widely known CLIP [66] model. Here we show results for a CLIP ViT-B student model that we finetuned on AligNet using the same setup and hyperparameters as we did for the SigLIP model. As can be seen in Fig.˜15, for few-shot learning the CLIP model generally benefits more from AligNet finetuning than SigLIP, while its final performance remains below that of the SigLIP model. The same pattern, can be seen for ImageNet and out-of-distribution performance (see Fig.˜13), and distribution shifts (see Fig.˜14). This further confirms the benefits of our method, and also justifies the focus on SigLIP in the main paper.
SigLIP2
The opposite pattern emerges when comparing SigLIP with the more recent SigLIP2 [81]: SigLIP2 has even stronger baseline performance than SigLIP and unsurprisingly benefits less from finetuning on AligNet. As can be seen in Fig.˜15 and Fig.˜14, it still does benefit in most cases for fewshot-generalization and distribution shift cases. For the out-of-distribution experiments in Fig.˜13, on the other hand, the effect is negligible or even detrimental. Given the strong baseline performance of SigLIP2 it would be fruitful to use the largest SigLIP2 as an even stronger teacher model, which we leave for future work.
B.2.5 Training on AligNet from Scratch
AligNet finetuning provides consistent improvements across many different models, settings, and datasets. This begs the question whether pretraining is actually needed, or if we can train a randomly initialized model solely on AligNet and achieve similar results. To answer this question, we trained a ViT-B/16 architecture from scratch on the AligNet dataset using the KLD triplet objective outlined in Eq.˜6. We optimize for 1M steps with the Adam optimizer using a cosine decay learning rate schedule with a peak learning rate of $0.0003$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay of $0.1$, $\tau^{\dagger}=1000$, and a batch size of $1024$. These hyperparameters were determined via grid search.
The results are shown in Figs.˜12, 13, 14 and 18, with the model listed as “Scratch”. They roughly fall into two clusters: The first one concerns the distribution-shift experiments on the Breeds dataset (see Fig.˜14), and odd-one-out accuracy on the THINGS dataset (see Fig.˜18a). Here our Scratch model performs almost as well as the other pretraining + finetuning models. This means that these tasks are quite closely related to the benefits that the AligNet fine-tuning provides.
The seconds cluster comprises the few-shot (see Fig.˜12) and ImageNet out-of-distribution results (see Fig.˜13). While the Scratch model improves substantially over training, on these tasks its performance clearly falls far short of that of the other pretrained models. For example it achieves approx. $27\%$ 1-shot accuracy on ImageNet compared to the $56\%$ accuracy of the same-sized SigLIP ViT-B model. Pretraining on other vision tasks is thus still important to augment AligNet training and clearly instills important complementary capabilities in a neural network model.
B.2.6 Perceptual patch similarity evaluations
Figure 16: Performance change on the BAPPS dataset after alignment. All changes are relatively small in absolute magnitude. Most of the AligNet models improve across both subsets of the benchmark, with the exception of the DINOv1 model, which shows significant improvements on Algorithms, but worsens on Distortions, and DINOv2, which changes nonsignificantly. These results suggest that AligNet tuning may slightly improve perceptual judgment performance in most cases. (Colors denote split subsets from the original work; transparent bars are non-signficant changes.)
Distortions Algorithms
Model cnn traditional color deblur frameinterp superres
CapPa-B base 84.62 59.26 63.62 60.83 54.87 65.33
CapPa-B AligNet 84.96 59.47 64.22 60.93 54.98 65.40
DINOv1-B base 85.23 60.28 63.79 60.96 55.08 65.47
DINOv1-B AligNet 84.87 59.36 65.59 60.97 55.72 65.65
DINOv2-B base 82.84 61.36 63.79 59.75 54.03 63.46
DINOv2-B AligNet 82.71 61.25 63.83 59.65 54.03 63.46
SigLIP-B base 84.34 59.39 62.78 61.01 54.18 65.10
SigLIP-B AligNet 84.79 59.75 63.69 60.95 54.45 65.13
ViT-B base 84.92 59.26 63.37 61.05 56.09 65.40
ViT-B AligNet 85.17 60.34 64.64 61.07 55.83 65.40
Table 7: Evaluations of models on the BAPP dataset, before and after AligNet tuning.
We also evaluated our AligNet models on the BAPPS Berkeley-Adobe Perceptual Patch Similarity (BAPPS) dataset [87]. This benchmark consists of human similarity judgments, but focuses on a different level of analysis—perceptual similarity of small image patches under distortion. Accordingly, we follow the original work in analyzing representations aggregated across the early layers of the models (specficially, we used the the first 5 transformer layers, analogous to the 5 convolutional layers used by the original authors). Analyses on this task therefore present a different test of how our alignment procedure affects model capacities—including the effect at earlier layers.
We show the changes between the base and AligNet models in Fig.˜16 and the absolute results in Tab.˜7. In general, the effects are very small—presumably both because this type of perceptual similarity is relatively different from the semantic structure we focused on, and because our alignment procedure is only explicitly targeted at representations much later in the model. Nevertheless, we generally see significant improvements (via exact binomial tests on the proportion of changed answers where AligNet improves over the base model) for the models across both subsets of the dataset. The exception is the DINO models — DINOv1 gets significantly worse at the Distortions subset, though it still improves significantly on Algorithms, while DINOv2 changes nonsignificantly for both. We find these results to be overall promising—they suggest that AligNet tuning is not dramatically harming other types of perceptual judgment performance, and in fact may slightly improve it for most models.
B.2.7 Ablations
@@
To investigate if AligNet finetuning affects models of different capacity/size in a similar way, we compared the performance of the ViT-B model (87M parameters) with the performances of a smaller ViT-S (22M parameters) and a larger ViT-L (305M parameters) model respectively. The results in Fig.˜12 and Fig.˜13 indicate that the positive impact of AligNet fine-tuning increases with model scale. For example, the average 10-shot performance improved by $2.4\%$ points for ViT-S, by $6.3\%$ points for ViT-B and by $8.3\%$ points for ViT-L. We observe similar results for all datasets we considered. Thus, we hypothesize that a larger model capacity makes it easier for a model to integrate the additional global coarse-grained information provided by AligNet’s human-like similarity judgments.
@@
We also wanted to see how AligNet scales with the amount of available training data. We individually varied the number of ImageNet images we used to construct our dataset over 3 orders of magnitude, and the number of triplets that we sampled across 4 orders of magnitude. We then looked at the 1-shot results across all the few-shot datasets described in Sec.˜B.2.1 (Outcomes for 10-shot were not qualitatively different). The results in Fig.˜17a show that AligNet performance only increases slightly as the number of available base images increases. However, as we can see from Fig.˜17b, the number of generated triplets does influence the outcome, though we hit diminishing returns eventually. We do not expect to be able to increase AligNet performance much further be simply increasing the amount of available data.
Figure 18: a: Odd-one-out accuracies on the THINGS dataset for different models in different settings. This figure is similar to (an extended version of) the THINGS plot in Fig.˜2A. “AligNet” and “Base Model” refer to the pretrained student model, without and without finetuning on AligNet respectively. UnAlignet refers to a student finetuned on a variant of the AligNet dataset with an unaligned teacher. b: Comparing 1-shot accuracy on 11 datasets after fine-tuning on UnaligNet ($x$-axis) with fine-tuning on AligNet ($y$-axis). Most points lie slightly above the diagonal indicating that alignment of the teacher leads to a small but significant improvement in 1-shot performance of the student. This ablation shows the importance of aligning the teacher model for getting aligned student models. c: A variation on b which colors each marker as green if it represents an improvement from AligNet finetuning, vs red for a degradation of performance. Note that 80 out of the 110 model/data combinations led to an improvement.
@@
An important part of the AligNet pipeline is to align the teacher model to human similarity judgments using a specially trained affine transformation. Here we investigate the effect of this step, by generating a variant of AligNet without this transformation that we call UnAligNet. For this dataset we do not use any human odd-one-out data, so finetuning on UnAligNet is equivalent to distilling the teacher model using our triplet sampling procedure. Fig.˜18A compares the odd-one-out accuracy of various student models when fine-tuning on AligNet vs on UnaligNet. As expected, we find that using an unaligned teacher leads to substantially worse alignment of the student models. Though it should be noted that some models (especially the ViTs trained on ImageNet) can already profit even from an unaligned teacher.
It is perhaps unsurprising that an aligned teacher improves the alignment of the student models. A more interesting question is, whether an aligned teacher also improves the students performance on other tasks such as few-shot classification. Fig.˜18B thus compares 1-shot performance after training on UnaligNet ($x$-axis) with performance after training on AligNet ($y$-axis). The first thing to note here, is that the differences between AligNet and UnaligNet are much smaller than the differences between AligNet and the Base Model (cf. Fig.˜12). This shows that the few-shot improvements are in large part due to the teacher model and our distillation method. But importantly, the vast majority of remaining improvements are positive, which means that using an aligned teacher gives a consistent advantage over an unaligned teacher.
@@
Fig.˜19 shows the 1-shot accuracy of student models when comparing different sampling strategies. In all three panels, the $y$-axis corresponds to the performance of a model after AligNet-fine-tuning. The left panel, again shows how this compares to the performance of the (not fine-tuned) base model ($x$-axis). The middle panel compares shows how AligNet performance compares to fine-tuning on random-triplets ($x$-axis). Finally, the right panel shows comparison between AligNet and fine-tuning on triplets sampled from class-borders. Notice that the difference in the middle panel between random triplet sampling is consistently and substantially worse than cluster-based sampling. This clearly highlights the importance of the triplet sampling for down-stream performance. The panel on the right, on the other hand, shows that there is virtually no difference between cluster-border sampling and class-border sampling. So when label information is available, that can effectively be used for sampling triplets, but even if there is not, the clustering based sampling will work just as well.
@@
We have stressed that fine-tuning models to better align them with human similarity judgments depends on having sufficiently large datasets, which do not exist yet. This argument has motivated our multi-stage approach of first linearly aligning a teacher model and then using this as a surrogate model to generate a large corpus of human-like triplet odd-one-out choices. Yet, if we look at Fig.˜17, it seems that decreasing the number of images or the number of triplets by up to two orders of magnitude still retains most of the benefit on downstream task performance (in terms of few-shot accuracy). This raises the question whether our pipeline as a whole might be unnecessary, given that the THINGS dataset [29] already contains 4.70 million human similarity judgments collected via online crowd- sourcing for 1854 object images.
Therefore, here we compare soft-alignment with using the THINGS odd-one-out triplets directly for fine-tuning student models. Fig.˜20 shows that fine-tuning directly on THINGS is consistently worse compared to training on AligNet in terms of ImageNet 1-shot accuracy ($x$-axis). Importantly, it appears that in almost all cases the performance of the model actually decreases below that of the base model. It is further interesting to note that even in terms of THINGS odd-one-out accuracy, the AligNet trained models perform better than the any other model. This is somewhat surprising, because it means that there clearly exists an overfitting problem when training on the THINGS data (even with strong $\ell_{2}$-regularization), which seems to mostly disappear when using AligNet for fine-tuning. One reason for this could be the differences in the sampling procedure (cluster-based vs. random sampling) and the loss function (soft-choices vs. hard-choices) between AligNet and THINGS.
@@
Uncertainty Distillation vs. gLocal. Here, we compare the Uncertainty Distillation (UD) transformation that we introduced in SI. A.1.1 against gLocal [cf. 56]. The learning objectives of the two linear transformations are highly similar and the goal is the same: the representations of a pretrained vision foundation model are linearly transformed into a global coarse-grained human object similarity space using the THINGS dataset [28] while the nearest neighbor structure of the model’s representation space is preserved. However, the difference between gLocal and UD is that UD directly injects the human uncertainty estimates for the triplet odd-one-out choices via approximate Bayesian inference [cf. 54] into a model’s representation space and, thus, uses soft rather than hard triplet choices (see SI. A.1.1 for details). We use the Levels datasets (see SI. A.3) for evaluating the UD and gLocal transformations, comparing the same set of models that were evaluated in Muttenthaler et al. [56]. The gLocal transformations for those models are publicly available via the Python library thingsvision [53] which we used for this analysis. In Fig.˜21 we see that the UD transformation is either equally accurate or better in aligning the model’s representations with the human similarity judgments, and predicting their disagreement levels and aggregated RTs respectively in the different abstraction level compared to gLocal. As before, human RTs are measured in log-space and model uncertainties are measured as discrete Shannon entropy over the triplet odd-one-out choices. Note that UD is computationally more efficient than gLocal because it does not require an extra set of images—in addition to the THINGS data—to evaluate the preservation of a representation’s nearest neighbor structure which the multi-objective of gLocal necessitates [cf. 56].